OpenStudy (anonymous):

The ratio of the width of the model windmill to the width of the real windmill provides the scale on which the city of Madurodam is built. The scale of the city is 1: ___. The real windmill width is 20 meters and the length is 30 meters. The model of the windmill is 1.2 meters and the length is 0.8 meters.

OpenStudy (anonymous):

This is a poorly phrased question. But we can figure out the answer by using the values that line up properly. 20/0.8 = 30/1.2 = scale But it is unclear whether they want that answer or 1/scale.